# HDCompression: Hybrid-Diffusion Image Compression for Ultra-Low Bitrates thanks: Equal Contribution

### v2

@@ Abstract
Image compression under ultra-low bitrates remains challenging for both conventional learned image compression (LIC) and generative vector-quantized (VQ) modeling. Conventional LIC suffers from severe artifacts due to heavy quantization, while generative VQ modeling gives poor fidelity due to the mismatch between learned generative priors and specific inputs. In this work, we propose Hybrid-Diffusion Image Compression (HDCompression), a dual-stream framework that utilizes both generative VQ-modeling and diffusion models, as well as conventional LIC, to achieve both high fidelity and high perceptual quality. Different from previous hybrid methods that directly use pre-trained LIC models to generate low-quality fidelity-preserving information from heavily quantized latent, we use diffusion models to extract high-quality complimentary fidelity information from the ground-truth input, which can enhance the system performance in several aspects: improving indices map prediction, enhancing the fidelity-preserving output of the LIC stream, and refining conditioned image reconstruction with VQ-latent correction. In addition, our diffusion model is based on a dense representative vector (DRV), which is lightweight with very simple sampling schedulers. Extensive experiments demonstrate that our HDCompression outperforms the previous conventional LIC, generative VQ-modeling, and hybrid frameworks in both quantitative metrics and qualitative visualization, providing balanced robust compression performance at ultra-low bitrates.
Index Terms:
Ultra-low Bitrate Image Compression, Diffusion Model.
I Introduction
@@
There has been an explosion of applications requiring transmitting large amounts of image data with limited bandwidth, calling for effective image compression solutions at ultra-low bitrates. Despite decades of research [1, 2, 3, 4], image compression at ultra-low bitrates remains an ongoing challenge. This is primarily due to the conventional framework of applying heavy quantization at ultra-low bitrates, resulting in significant artifacts. In the conventional framework, an encoder first transforms the input image into a latent feature, either by traditional transformation as in JPEG [1] or by neural network models as in learned image compression (LIC) [4]. Then the latent feature is quantized by rounding operations for transmission. A decoder subsequently recovers the output image from the dequantized latent feature using traditional inverse-transformation or neural network models. Therefore, bit reduction occurs during the quantization process. Especially at ultra-low bitrates, intense quantization causes excessive information loss, leading to severe and unpleasant blurriness and noises. Although numerous efforts [4, 5, 6] have been focused on improving the transformation model and prediction of quantization statistics, such artifacts cannot be easily mitigated, as shown in Fig. 1.
@@
Besides Variational AudoEncoder (VAE), generative methods such as generative adversarial networks (GANs) [7, 8, 9, 10] and diffusion models [11, 12, 13, 14, 15] offer promising opportunities to explore alternative frameworks for image compression. Generative approaches learn statistical priors from images, which allows for the synthesis of perceptually realistic high-quality image details from degraded input images. For instance, vector-quantized (VQ) image modeling [16, 17] has been recently used for image compression [18, 19, 20], where the learned generative priors serve as visual codewords that span a latent space, enabling images to be mapped into vector-quantized integer indices. Thus, the learned VQ latent space provides a refined quantization strategy that retrieves high-quality, information-rich codewords for reconstructing high-realism outputs, which could lead to finer quantization adjustments and effectively avoid the degraded outputs at ultra-low bitrates. However, while the generated outputs are visually appealing to human eyes, the learned generative priors (i.e., codewords) often deviate from authentic image details, bringing about significant pixel-level differences from the original inputs. Thus, most VQ-modeling-based approaches [18, 19] primarily address perceptual quality only and operate at extremely low bitrates where poor fidelity may be tolerated, as shown in Fig. 1.
@@
For the practical task of image compression, both content authenticity and visual quality are crucial, even at ultra-low bitrates. However, there is a complex and contradictory relationship between perceptual quality and fidelity [21], making it very challenging for a method to perfect both aspects for general scenarios. Recently, HybridFlow [22] has synergized the conventional LIC and the generative VQ-modeling to preserve both fidelity and perceptual quality at ultra-low bitrates (around 0.05 bpp). In HybridFlow, VQ-modeling provides high-realism generation, while conventional LIC offers authentic details from each specific input. However, after incorporating conventional LIC directly into HybridFlow, the quantization issue at ultra-low bitrates still severely impacts the assistive fidelity information quality for indices map prediction and conditional reconstruction.
@@
To address the issues above and maintain the balance between fidelity and perceptual quality, we propose the Hybrid-Diffusion Compression (HDCompression) approach that effectively exploits both GAN and diffusion models (DM) for ultra-low-bitrate image compression in a dual manner. The diffusion model contributes detailed perceptual features that address high-fidelity requirements, while the VQ-based stream provides structured latent codebooks that enable efficient compression at ultra-low bitrates. Furthermore, instead of directly using pre-trained DM as previous DM-based compression methods [23], we utilize a dense representative vector (DRV) to mitigate the heavy computation and memory consumption issues. Compared with previous state-of-the-art (SOTA) image compression methods, including traditional VVC, single-stream LIC method MLIC [6], single-stream VQ-modeling method [19], and dual-stream HybridFlow [22], HDCompression can improve perceptual quality (LPIPS) by 26% over HybridFlow, while maintaining the same level of fidelity (PSNR), and can reduce artifacts like random or structured noise patterns. We highlight our contributions as follows.
@@
II Related Work
II-A Learned Image Compression
LIC [4, 6, 24] using neural networks [25, 26, 27] has shown superior performance over traditional methods like JPEG [1], VVC [2]. One most popular LIC frameworks is based on VAE. In the encoder, the input image is encoded into a dense latent feature, which is quantized by rounding operations for efficient transmission. Then the decoder reconstructs the output image based on the dequantized latent feature. At ultra-low bitrates, the information loss caused by the universal rounding quantization is too damaging to be recovered by the decoder, resulting in severe artifacts like blurriness, noises, blocky effects, etc.
II-B Image Compression by Generative Models
@@
VQ-codebook-based image compression. GAN-assisted VQ-codebook methods have shown advantages over traditional LICs, especially in perceptual quality [20, 19, 28]. A discrete codebook maps the encoded image latent into a transmitted integer indices map instead of the conventional rounding-quantized latent, enabling lower bitrate and greater resilience to network fluctuations. The decoder retrieves the vector-quantized (VQ) latent from the shared codebook according to the indices map and then reconstructs the image. To improve the visual quality and fidelity of the reconstruction from the VQ-based generator, the GAN discriminator is either applied on the pixel level [20, 19] or further extended to the indices map [28]. However, the VQ-codebook captures general image priors and often deviates from individual image details. As a result, although visually appealing, the reconstructed images usually present significant pixel-level distortions (e.g., poor PSNR).
@@
Image compression with diffusion models. DMs have surpassed GAN in many vision tasks. Based on VAE and latent diffusion, latent diffusion models (LDMs) can be directly applied to image compression with minor changes. For example, hyperpriors extracted from the input are used as conditional information for the multi-step denoising process to recover a denoised latent for reconstruction [29]. The performance suffers at ultra-low bitrates as the noisy initialization requires relatively accurate denoising guidance. Also, many steps ($>$15) are usually required, causing severe computation latency. The problem may be alleviated by exploiting the strong generative ability of pre-trained LDM like Stable Diffusion (SD), where the rounding-quantized latent is refined by inverse diffusion [23] with adaptive denoising steps. However, the severe memory burden of pre-trained SD models (often with $>$1.3B parameters) is impractical for compression.
@@
Hybrid dual-stream image compression. A dual-stream LIC framework has been proposed recently, which combines the VQ-codebook-based compression and conventional LIC and takes advantage of the resulting synergy at ultra-low bitrates. For example, HybridFlow [22] uses pre-trained conventional LIC models to provide fidelity information, which assists the VQ-codebook-based stream in both indices prediction and generative reconstruction. The dual-stream framework aims to achieve a balanced reconstruction quality between fidelity and perception at ultra-low bitrates. However, conventional LIC methods are directly equipped into the entire system without any adaptation. The quality of the assisting fidelity information still suffers from the common rounding quantization issue of general LIC. Inspired by the dual-stream framework, we use DMs in this paper to effectively provide complementary input-specific fidelity information to boost performance further.
II-C Dense-Vector-based Vision Model
@@
For convolution-based vision models [30, 31, 32], an input image is commonly encoded into a latent feature $L$ as a 3D tensor. As transformer blocks gain popularity in vision models, it has been shown that a highly dense 1D vector $V$ is quite powerful to serve as conditional guidance for various downstream tasks. In general, $V$ carries input-adaptive information learned for specific tasks, which conditionally modify $L$ to improve performance over individual inputs. For instance, RCG [33] uses a global guidance $V$ as a condition for image generation. DiffIR [34] uses a joint embedding $V$ between the ground-truth and degraded inputs to provide ground-truth information for guiding image restoration. In this work, we incorporate such a dense vector $V$ to provide complementary input-specific fidelity information for improved reconstruction.
III Methodology
@@
III-A The LIC Stream
For general LIC, an input image $\textbf{x}\in\mathbb{R}^{H\!\times\!W\!\times\!3}$ is first encoded into a latent $\textbf{y}\in\mathbb{R}^{\frac{H}{m}\!\times\!\frac{W}{m}\!\times\!c}$ by an LIC encoder. Then y is rounding quantized into $\textbf{y}_{q}$ for easy transmission, and a LIC decoder reconstructs the output image ${\hat{\textbf{x}}}$ from received $\textbf{y}_{q}$. The downsampling factor $m$ and latent dimension $c$ determine y’s size, e.g., a larger y that has more representative capacity, gives better reconstruction but consumes more bits.
In the dual-stream framework, the LIC stream provides the fidelity information to the final reconstruction. At ultra-low bitrates, such information quality severely suffers due to the large rounding loss. In this work, instead of merely relying on pre-trained LIC models, we introduce a DRV-diffusion-based enhancement module to improve the fidelity of information from the LIC stream. Following the DRV-based vision model, we leverage a DRV to carry ground-truth information from the current input x to enhance the fidelity of the LIC stream. In detail, the structure of the DRV-based enhancement module is inspired by DiffIR [34], where a DRV $\textbf{v}_{gt}$ containing ground-truth information is fused into Restormer via cross-attention in transformer blocks. However, $\textbf{v}_{gt}$ is too heavy to transfer for ultra-low-bitrate compression. Therefore, we propose to regenerate a DRV in the decoder by utilizing a joint-embedding DRV vector $\textbf{v}_{joint_{E}}$:
$\textbf{v}_{joint_{E}}=E_{L}(\textbf{x},\hat{\textbf{x}}),$ (1)
@@ III-A The LIC Stream
where $t\in[0,T]$ is the iterative denoising process, $\alpha_{t}=1-\beta_{t}$. Vector c has the same shape as $\textbf{v}_{joint_{E}}$, and is extracted from ${\hat{\textbf{x}}}$ by a DRV extractor $E_{C}$ that is forked from $E_{L}$ but has a modified input dimension to take only ${\hat{\textbf{x}}}$ as input. Since our DRV only needs to provide complementary fidelity information instead of modeling entire image structures, a simple denoising network $\epsilon_{\theta}$ consisting of 4 ResMLP blocks with a 4-step sampling scheduler is used, largely reducing computation and memory requirements compared with conventional UNet-based SD denoising. $\hat{\textbf{v}}_{joint_{E}}$ is then embedded into the Restormer [35] for latent enhancement via additional cross-attention blocks inserted into the inner transformer blocks where $\hat{\textbf{v}}_{joint_{E}}$ serves as $key$ and $value$. Finally, the enhanced latent is fed into the LIC decoder to generate the final output image ${\hat{\textbf{x}}}_{lic}$ from the LIC stream, serving as the interactive baseline with the codebook-based generative stream for fidelity infusion.
III-B The Generative Stream
@@
Transmitting whole $\textbf{d}_{vq}$ can be too costly for ultra-low-bitrate scenarios. For example for $K$=1024 and $n$=16, the bpp is about $10/256\!\!\approx\!\!0.04$, which is close to the upper bound of ultra-low bitrate range ($<$0.05). Thus, compression-friendly binary masks have been used [22] to transmit only a masked portion of $\textbf{d}_{vq}$: $\textbf{d}_{masked}=\textbf{m}\circ\textbf{d}_{vq}$.
@@
Figure 3: DRV $\hat{\textbf{v}}_{joint_{P}}$ embedding process in T.
III-B1 LIC-assisted indices map prediction
The token-based encoder-decoder transformer T takes as input a token-embedding $\textbf{emb}_{P}$ with length of $\frac{HW}{n^{2}}\!+\!1$ by $\{emb_{token_{0}},emb_{id_{0}},emb_{mask},emb_{id_{3}},\cdots,emb_{id_{\frac{HW}{n^{2}}}}\}$, where $mask$ is the masked index to predict and $emb_{mask}$ is the embedding vector of $mask$; $id_{i}$ is the unmasked ground-truth index and $emb_{id_{i}}$ is its embedding vector; and $token_{0}$ is a class token originally designed for class-based generation and $emb_{token_{0}}$ is the embedding vector of $token_{0}$. HybridFlow [22] ignores $token_{0}$ by using a fake class label without actual meaning. However, as shown in MAGE [36], instead of fake ‘dummy’ embedding, global priors can be fused into masked locations for improved prediction. Therefore, we fuse information from ${\hat{\textbf{x}}}$ into $\textbf{emb}_{P}$, which serves as image-aware global embedding for better prediction.

### v3

@@ Abstract
Image compression under ultra-low bitrates remains challenging for both conventional learned image compression (LIC) and generative vector-quantized (VQ) modeling. Conventional LIC suffers from severe artifacts due to heavy quantization, while generative VQ modeling gives poor fidelity due to the mismatch between learned generative priors and specific inputs. In this work, we propose Hybrid-Diffusion Image Compression (HDCompression), a dual-stream framework that utilizes both generative VQ-modeling and diffusion models, as well as conventional LIC, to achieve both high fidelity and high perceptual quality. Different from previous hybrid methods that directly use pre-trained LIC models to generate low-quality fidelity-preserving information from heavily quantized latent, we use diffusion models to extract high-quality complementary fidelity information from the ground-truth input, which can enhance the system performance in several aspects: improving index map prediction, enhancing the fidelity-preserving output of the LIC stream, and refining conditioned image reconstruction with VQ-latent correction. In addition, our diffusion model is based on a dense representative vector (DRV), which is lightweight with very simple sampling schedulers. Extensive experiments demonstrate that our HDCompression outperforms the previous conventional LIC, generative VQ-modeling, and hybrid frameworks in both quantitative metrics and qualitative visualization, providing balanced robust compression performance at ultra-low bitrates.
footnotetext: *Both authors contributed equally.
1 Introduction
@@
There has been an explosion of applications requiring transmitting large amounts of image data with limited bandwidth, calling for effective image compression solutions at ultra-low bitrates. Despite decades of research [jpegai, BPG, vvc, JPEG2000], image compression at ultra-low bitrates remains an ongoing challenge. This is primarily due to the conventional framework of applying heavy quantization at ultra-low bitrates, resulting in significant artifacts. In the conventional framework, an encoder first transforms the input image into a latent feature, either by traditional transformation as in JPEG [JPEG2000] or by neural network models as in learned image compression (LIC) [jpegai]. Then the latent feature is quantized by rounding operations for transmission. A decoder subsequently recovers the output image from the dequantized latent feature using traditional inverse-transformation or neural network models. Therefore, bit reduction occurs during the quantization process. Especially at ultra-low bitrates, intense quantization causes excessive information loss, leading to severe and unpleasant blurriness and noises. Although numerous efforts [jpegai, losslesshyperprior, mlic2023] have been focused on improving the transformation model and prediction of quantization statistics, such artifacts cannot be easily mitigated, as shown in Fig. 1.
@@
Besides Variational AudoEncoder (VAE), generative methods such as generative adversarial networks (GANs) [goodfellow2014generative, isola2017image, mirza2014conditional, radford2015unsupervised] and diffusion models [dhariwal2021diffusion, ho2020denoising, kong2025, li2024pruning, rombach2022high, zeng2025enhancing] offer promising opportunities to explore alternative frameworks for image compression. Generative approaches learn statistical priors from images, which allows for the synthesis of perceptually realistic high-quality image details from degraded input images. For instance, vector-quantized (VQ) image modeling [esser2021taming, van2017neural] has been recently used for image compression [jia2024generative, WACV2024, vqganmasiwei], where the learned generative priors serve as visual codewords that span a latent space, enabling images to be mapped into vector-quantized integer indices. Thus, the learned VQ latent space provides a refined quantization strategy that retrieves high-quality, information-rich codewords for reconstructing high-realism outputs, which could lead to finer quantization adjustments and effectively avoid the degraded outputs at ultra-low bitrates. However, while the generated outputs are visually appealing to human eyes, the learned generative priors (i.e., codewords) often deviate from authentic image details, bringing about significant pixel-level differences from the original inputs. Thus, most VQ-modeling-based approaches [jia2024generative, vqganmasiwei] primarily address perceptual quality only and operate at extremely low bitrates where poor fidelity may be tolerated, as shown in Fig. 1.
@@
For the practical task of image compression, both content authenticity and visual quality are crucial, even at ultra-low bitrates. However, there is a complex and contradictory relationship between perceptual quality and fidelity [blau2018perception], making it very challenging for a method to perfect both aspects for general scenarios. Recently, HybridFlow [lu2024hybridflow] has synergized the conventional LIC and the generative VQ-modeling to preserve both fidelity and perceptual quality at ultra-low bitrates (around 0.05 bpp). In HybridFlow, VQ-modeling provides high-realism generation, while conventional LIC offers authentic details from each specific input. However, after incorporating conventional LIC directly into HybridFlow, the quantization issue at ultra-low bitrates still severely impacts the assistive fidelity information quality for indices map prediction and conditional reconstruction.
@@
To address the issues above and maintain the balance between fidelity and perceptual quality, we propose the Hybrid-Diffusion Compression (HDCompression) approach that effectively exploits both GAN and diffusion models (DM) for ultra-low-bitrate image compression in a dual manner. The diffusion model contributes detailed perceptual features that address high-fidelity requirements, while the VQ-based stream provides structured latent codebooks that enable efficient compression at ultra-low bitrates. Furthermore, instead of directly using pre-trained DM as previous DM-based compression methods [relic2024lossy], we utilize a dense representative vector (DRV) to mitigate the heavy computation and memory consumption issues. Compared with previous state-of-the-art (SOTA) image compression methods, including traditional VVC, single-stream LIC method MLIC [mlic2023], single-stream VQ-modeling method [vqganmasiwei], and dual-stream HybridFlow [lu2024hybridflow], HDCompression can improve perceptual quality (LPIPS) by 26% over HybridFlow, while maintaining the same level of fidelity (PSNR), and can reduce artifacts like random or structured noise patterns. We highlight our contributions as follows.
@@
2 Related Work
2.1 Learned Image Compression
LIC [jpegai, Cheng2020, mlic2023] using neural networks [Hyperprior2018, li2024nas, ma2019image] has shown superior performance over traditional methods like JPEG [JPEG2000], VVC [vvc]. One most popular LIC frameworks is based on VAE. In the encoder, the input image is encoded into a dense latent feature, which is quantized by rounding operations for efficient transmission. Then the decoder reconstructs the output image based on the dequantized latent feature. At ultra-low bitrates, the information loss caused by the universal rounding quantization is too damaging to be recovered by the decoder, resulting in severe artifacts like blurriness, noises, blocky effects, etc.
2.2 Image Compression by Generative Models
@@
VQ-codebook-based image compression. GAN-assisted VQ-codebook methods have shown advantages over traditional LICs, especially in perceptual quality [WACV2024, vqganmasiwei, muckley2023improving]. A discrete codebook maps the encoded image latent into a transmitted integer indices map instead of the conventional rounding-quantized latent, enabling lower bitrate and greater resilience to network fluctuations. The decoder retrieves the vector-quantized (VQ) latent from the shared codebook according to the indices map and then reconstructs the image. To improve the visual quality and fidelity of the reconstruction from the VQ-based generator, the GAN discriminator is either applied on the pixel level [WACV2024, vqganmasiwei] or further extended to the indices map [muckley2023improving]. However, the VQ-codebook captures general image priors and often deviates from individual image details. As a result, although visually appealing, the reconstructed images usually present significant pixel-level distortions (e.g., poor PSNR).
@@
Image compression with diffusion models. DMs have surpassed GAN in many vision tasks. Based on VAE and latent diffusion, latent diffusion models (LDMs) can be directly applied to image compression with minor changes. For example, hyperpriors extracted from the input are used as conditional information for the multi-step denoising process to recover a denoised latent for reconstruction [yang2024lossy]. The performance suffers at ultra-low bitrates as the noisy initialization requires relatively accurate denoising guidance. Also, many steps (>>15) are usually required, causing severe computation latency. The problem may be alleviated by exploiting the strong generative ability of pre-trained LDM like Stable Diffusion (SD), where the rounding-quantized latent is refined by inverse diffusion [relic2024lossy] with adaptive denoising steps. However, the severe memory burden of pre-trained SD models (often with >>1.3B parameters) is impractical for compression.
@@
Hybrid dual-stream image compression. A dual-stream LIC framework has been proposed recently, which combines the VQ-codebook-based compression and conventional LIC and takes advantage of the resulting synergy at ultra-low bitrates. For example, HybridFlow [lu2024hybridflow] uses pre-trained conventional LIC models to provide fidelity information, which assists the VQ-codebook-based stream in both indices prediction and generative reconstruction. The dual-stream framework aims to achieve a balanced reconstruction quality between fidelity and perception at ultra-low bitrates. However, conventional LIC methods are directly equipped into the entire system without any adaptation. The quality of the assisting fidelity information still suffers from the common rounding quantization issue of general LIC. Inspired by the dual-stream framework, we use DMs in this paper to effectively provide complementary input-specific fidelity information to boost performance further.
2.3 Dense-Vector-based Vision Model
@@
For convolution-based vision models [cavigelli2017cas, li2023less, li2025achieving, li2025frequency], an input image is commonly encoded into a latent feature $L$ as a 3D tensor [gong2022reverse, li2025sepprune, li2024sglp, yang2025fairsmoe]. As transformer blocks gain popularity in vision models, it has been shown that a highly dense 1D vector $V$ is quite powerful to serve as conditional guidance for various downstream tasks. In general, $V$ carries input-adaptive information learned for specific tasks, which conditionally modify $L$ to improve performance over individual inputs. For instance, RCG [RCG2023] uses a global guidance $V$ as a condition for image generation. DiffIR [xia2023diffir] uses a joint embedding $V$ between the ground-truth and degraded inputs to provide ground-truth information for guiding image restoration. In this work, we incorporate such a dense vector $V$ to provide complementary input-specific fidelity information for improved reconstruction.
3 Methodology
@@
3.1 The LIC Stream
For general LIC, an input image $\textbf{x}\in\mathbb{R}^{H\!\times\!W\!\times\!3}$ is first encoded into a latent $\textbf{y}\in\mathbb{R}^{\frac{H}{m}\!\times\!\frac{W}{m}\!\times\!c}$ by an LIC encoder. Then y is rounding quantized into $\textbf{y}_{q}$ for easy transmission, and a LIC decoder reconstructs the output image $^{\hat{\textbf{x}}}$ from received $\textbf{y}_{q}$. The downsampling factor $m$ and latent dimension $c$ determine y’s size, e.g., a larger y that has more representative capacity, gives better reconstruction but consumes more bits.
In the dual-stream framework, the LIC stream provides the fidelity information to the final reconstruction. At ultra-low bitrates, such information quality severely suffers due to the large rounding loss. In this work, instead of merely relying on pre-trained LIC models, we introduce a DRV-diffusion-based enhancement module to improve the fidelity of information from the LIC stream. Following the DRV-based vision model, we leverage a DRV to carry ground-truth information from the current input x to enhance the fidelity of the LIC stream. In detail, the structure of the DRV-based enhancement module is inspired by DiffIR [xia2023diffir], where a DRV $\textbf{v}_{gt}$ containing ground-truth information is fused into Restormer via cross-attention in transformer blocks. However, $\textbf{v}_{gt}$ is too heavy to transfer for ultra-low-bitrate compression. Therefore, we propose to regenerate a DRV in the decoder by utilizing a joint-embedding DRV vector $\textbf{v}_{joint_{E}}$:
[…]
@@ 3.1 The LIC Stream
where $t\in[0,T]$ is the iterative denoising process, $\alpha_{t}=1-\beta_{t}$. Vector c has the same shape as $\textbf{v}_{joint_{E}}$, and is extracted from $^{\hat{\textbf{x}}}$ by a DRV extractor $E_{C}$ that is forked from $E_{L}$ but has a modified input dimension to take only $^{\hat{\textbf{x}}}$ as input. Since our DRV only needs to provide complementary fidelity information instead of modeling entire image structures, a simple denoising network $\epsilon_{\theta}$ consisting of 4 ResMLP blocks with a 4-step sampling scheduler is used, largely reducing computation and memory requirements compared with conventional UNet-based SD denoising. $\hat{\textbf{v}}_{joint_{E}}$ is then embedded into the Restormer [zamir2022restormer] for latent enhancement via additional cross-attention blocks inserted into the inner transformer blocks where $\hat{\textbf{v}}_{joint_{E}}$ serves as keykey and valuevalue. Finally, the enhanced latent is fed into the LIC decoder to generate the final output image ${\hat{\textbf{x}}}_{lic}$ from the LIC stream, serving as the interactive baseline with the codebook-based generative stream for fidelity infusion.
Figure 3: DRV $\hat{\textbf{v}}_{joint_{P}}$ embedding process in T.
3.2 The Generative Stream
@@
Transmitting whole $\textbf{d}_{vq}$ can be too costly for ultra-low-bitrate scenarios. For example for KK=1024 and nn=16, the bpp is about $10/256\!\!\approx\!\!0.04$, which is close to the upper bound of ultra-low bitrate range (<<0.05). Thus, compression-friendly binary masks have been used [lu2024hybridflow] to transmit only a masked portion of $\textbf{d}_{vq}$: $\textbf{d}_{masked}=\textbf{m}\circ\textbf{d}_{vq}$.
@@
3.2.1 LIC-assisted indices map prediction
The token-based encoder-decoder transformer T takes as input a token-embedding $\textbf{emb}_{P}$ with length of $\frac{HW}{n^{2}}\!+\!1$ by $\{emb_{token_{0}},\\ emb_{id_{0}},emb_{mask},emb_{id_{3}},\cdots,emb_{id_{\frac{HW}{n^{2}}}}\}$, where maskmask is the masked index to predict and $emb_{mask}$ is the embedding vector of maskmask; $id_{i}$ is the unmasked ground-truth index and $emb_{id_{i}}$ is its embedding vector; and $token_{0}$ is a class token originally designed for class-based generation and $emb_{token_{0}}$ is the embedding vector of $token_{0}$. HybridFlow [lu2024hybridflow] ignores $token_{0}$ by using a fake class label without actual meaning. However, as shown in MAGE [li2023mage], instead of fake ‘dummy’ embedding, global priors can be fused into masked locations for improved prediction. Therefore, we fuse information from $^{\hat{\textbf{x}}}$ into $\textbf{emb}_{P}$, which serves as image-aware global embedding for better prediction.